OpenStudy (anonymous):

A construction company charges $15 per hour for debris removal, plus a one-time fee for the use of a trash dumpster. The total fee for 9 hours of service is $195. a.) Write the point slope form of an equation to find the total fee "y" for any number of hours "x" b.) Write the equation in slop-intercept form. c.) What is the fee for the use of a trash dumpster?

OpenStudy (anonymous):

a) y-195=15(x-9) b) y=15x+b 195=15(9)+b b=60 thus,y=15x+60 c)$60

OpenStudy (anonymous):

Thanks(: